Problem 4 [20 points]. The proportion of contaminant X and Y found among the con- taminants of a soil sample have the followic) (5 points) Determine whether the presence of contaminant X is independent of the presence of contaminant Y. Explain your ce) (4 points) On average, what proportion of contaminant in a soil sample is made up of Y, if 50% is contaminant X?
Answer #1

Solution:

The joint PDF is f\left ( x,y \right )=cxy^2;0\leqslant x+y\leqslant 1 .

a)The support of the PDF is shown shaded below.

x+y=1

b) The condition for PDF is

f(x,y) = 1 cxy2dydx = 1 0 J0 0 Jo x (1-x)dx = 1 x3 (1-x)dx=1 CX 0 .1 .1 (1.4-1/5) = 1

60 = 1 с 60

c) The marginal PDFs are

fx (x) - f (x, y) dy 0 fx (X) cxy2dy 0 f(x, y) dx 0 cxy2dx 0

Since f\left ( x,y \right )\neq f_X\left ( x \right )f_Y\left ( y \right ) , X,Y are not independent.

d) The conditional PDF,

(x.y 60xy (y|x) = 20x(1-X) зу fyxo(y|x) = (1-x

When x=0.5,

зуг fyX= 0.5 (y|x) (1-0.5)

The conditional probability,

P\left ( Y\geqslant 0.25|X=0.5 \right )=\int_{0.25}^{0.5}f_{Y|X=0.5}\left ( y|x \right )dy\\ P\left ( Y\geqslant 0.25|X=0.5 \right )=\int_{0.25}^{0.5}\frac{3y^2}{\left ( 1-0.5 \right )^3}dy\\ P\left ( Y\geqslant 0.25|X=0.5 \right )=\frac{\left [ y^3 \right ]_{0.25}^{0.5}}{\left ( 1-0.5 \right )^3}\\ P\left ( Y\geqslant 0.25|X=0.5 \right )=\frac{0.125-0.25^3}{0.125}\\ {\color{Blue} P\left ( Y\geqslant 0.25|X=0.5 \right )=0.875}